OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 5x − 2x = 2x + 2 5x + 2 = 2x − 10 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describes solutions to equations shown.

OpenStudy (anonymous):

Equation #1 5x-2x=2x+2 3x=2x+2 x=2 Equation #2 5x+2=2x-10 3x=-12 x=-4 The solution to equation #1 is larger than the solution to equation #2.